OpenStudy (ujjwal):

A fish rising vertically to the surface of water in a lake uniformly at the rate of 3m/s observes a kingfisher diving vertically towards the water at a rate of 9m/s vertically above it. If the R.I. of water is 4/3, find the actual velocity of dive of bird.

OpenStudy (experimentx):

looks like question is related to "related rates"

OpenStudy (ujjwal):

And refraction.. + a lot of logic.. I guess..

OpenStudy (vincent-lyon.fr):

Is the answer 8 m/s ? or 4.5 m/s ? I am not so sure of the relation for the plane boundary.

OpenStudy (ujjwal):

its 4.5 m/s

OpenStudy (ujjwal):

i got 8 m/s

OpenStudy (experimentx):

looks like the actual height is less than observed height!!

OpenStudy (vincent-lyon.fr):

Velocity of virtual bird is 9-3 = 6 m/s Seen from underwater, the virtual bird appears further away than actual bird by a factor n'/n = 4/3. So actual velocity is 6 / (4/3) = 6 x 3/4 = 4.5 m/s

OpenStudy (ujjwal):

yes, since it is viewed from denser to rarer medium @experimentX

OpenStudy (ujjwal):

Thanks @Vincent-Lyon.Fr and @experimentX